You have 50.0 mL of pure water and in this you dissolve 24.0 g of sugar. You then add this to 80 mL salt water with a density of 1.25 g/cm3. What is the overall density of the sugar-salt solution?

  1. What is the mass of the salt solution? What is the mass of the pure water? What is the mass of the sugar solution?
  2. What is the mass of the sugar-salt solution?
  3. What is the volume of the salt solution? What is the volume of the pure water? What is the volume of the sugar solution?
  4. What is the volume of the sugar-salt solution?
1
Expert's answer
2021-03-23T07:25:34-0400

1.the mass of the salt solution:

"m=\rho\times V=1.25\times80=100" g

the mass of the pure water:

"m=\rho\times V=1.00\times50=50" g

the mass of the sugar solution:

50+24=74

2.The mass of the sugar-salt solution

74+100=174

3.the volume of the salt solution:

"V=80"

the volume of the pure water

"V=50"

the volume of the sugar solution:

"V=\frac{m}{\rho}=\frac{74}{1.587}=46.63"

4.What is the volume of the sugar-salt solution

46.63+80=126.63


density of the sugar-salt solution:

"\rho=\frac{m}{V}=\frac{174}{126.63}=1.37"